Shift Summary for owl shift, 07/02/99

Sandrine Laplace (laplace@hep.uchicago.edu)
Fri, 2 Jul 1999 06:54:35 -0500

Shift Summary for owl shift, 07/02/99

On Shift: Sandrine, Jason L.

Hours of beam delivered at all intensities: 6.75
Hours of beam delivered at nominal intensity: 6.75
Hours of E832 data (special runs not included): 2.50
Hours of special runs (laser scans, muon runs...): 0.00

Accounting of lost time:

-- ETOT problem. Made an access of more than 2 hours !
-- HCC DYC full.

========================================================================
Runs ended this shift:
13788 - first 23 spills fine, then beam problems, and ETOT problems.
13789, 13790 - runs to test ETOT and HCC.

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

Sydney, Jason, Sandrine, Ed went on access for 2.5 hours.

========================================================================
Overview of shift:

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

00:00 It was really really hard to get up to go
to this shift...It's gonna be long...

13788 00:01 Arrive when evening shifties switch to muon
run. Then begin e832 run number 13788.

13788 00:45 The beam is quite irregular. MCR is often
aborting.

13788 00:57 Beam is gone for 15 min. now.
ETOT failure : low voltage of Bottom West is down.
Jason H. and Sydney paged.

13788 1:30 Try to cycle the Voltage supplies upstairs.
Doesn't help. Have to make an access.
Ed paged.

13788 1:45 Jason arrives. He and Sandrine make an access.
Swap the strip for these power supplies.
Doesn't fix it.

13788 1:49 End of run 13788.

2:21 Sydney joins on access with Jason and Sandrine.
A fuse of one of the power supply is down.
Swapping
it doesn't help, this PS still doesn't work.
Swap the PS itself.

3:15 Ed joins on access.
After many failures, the PSs finally work again.

13789 3:55 Start run (without beam) 13889, then 13890, to
test ETOT.
Another problem with ETOT appears concerning the
temperatures of boards 41 -> 45.
HCC DYC full. Never fixed by kfrend.

13890 4:15 Ask for beam. HCC still doesn't work.

13890 4:20 Beam is back. Disable triggers with HCC, we get
triggers. But lose permit many times.

4:55 Swap HCC controller. Everything's finally fine !

13891 5:00 Start e832 run 13891.

5:10 Can finally do the great job of a shifty : look
at online histos, while everybody who's not on
shift
go to have waffles and pancakes. Asked for a
waffle
with strawberries and cream to go, but wouldn't
bet on it...

5:30 I just saw half of the sun behin the landscape.
This is the moment I prefer on owl shift.
And yes, I'm still looking at online histos....

5:35 Gas check list completed.

06:20 Notice that the Munters has tripped off, thus
explaining the unusual warmth of the control room.
The day shift should get a hold of Leonard Nelson
about this.

7:05 shift is over, I can actually finally enjoy my
waffle....

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

********************************************************************
* LAPLACE Sandrine *
* --------------- *
* Enrico Fermi Institute KTeV 5514,S Blackstone Apt 318*
* 5640, S Ellis ave. at Chicago, IL 60637 *
* Chicago, IL 60637, USA Fermilab USA *
* http://www.geocities.com/Nashville/Rodeo/1190 *
********************************************************************